OpenStudy (anonymous):

What is the perimeter of a square that has a side of \sqrt{72} ? Write your answer in simplest radical form.

OpenStudy (zzr0ck3r):

perimeter of a square is x+x+x+x = 4x where x is the length of the side

OpenStudy (anonymous):

i dont think there is a square root of 72

OpenStudy (anonymous):

there is

OpenStudy (zzr0ck3r):

sure there is. its just not "Nice"

OpenStudy (zzr0ck3r):

you can take the square root of any non negative real number

OpenStudy (anonymous):

\[24\sqrt{4}\]

OpenStudy (mathstudent55):

The perimeter, P, of a square is 4 times the side, s. \(P = 4s\) \(P = 4 \times \sqrt{72} \) Now you need to simplify.

OpenStudy (mathstudent55):

sqrt(4) = 2

OpenStudy (zzr0ck3r):

um sqrt(72) = sqrt(36*2) = sqrt(36)*sqrt(2) = 6swrt(2)

OpenStudy (zzr0ck3r):

square root of any prime is irrational, thus we just leave it alone. because there is NO other way to write it.

OpenStudy (anonymous):

uhh dont u have to do something to make it 4 yimrd

OpenStudy (zzr0ck3r):

yimrd?

OpenStudy (anonymous):

times i meant to write times

OpenStudy (anonymous):

Im so confused.

OpenStudy (mathstudent55):

\(P = 4 \times \sqrt{72} = 4 \times \sqrt{36 \times 2} =4 \times 6 \sqrt{2} = 24 \sqrt{2} \)

OpenStudy (zzr0ck3r):

you have a square with side sqrt(72), so the perimeter is 4*sqrt(72) sqrt(72) is just a number..... if your side was 10 then it would be 10*4

OpenStudy (anonymous):

so what would be the answer

OpenStudy (anonymous):

the answer would be \[24\sqrt{2}\]

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

4*6*sqrt(2) = 4*sqrt(72)